# Adaptive Gaussian Mixture Models-Based Anomaly Detection for Under-Constrained Cable-Driven Parallel Robots

### v1

@@ Abstract
keywords:
Anomaly detection, Cable-Driven Parallel Robots, Gaussian Mixture Models, Unsupervised learning, Robotics control systems
Julio Garrido* Javier Vales Diego Silva-Muñiz Enrique Riveiro Pablo López-Matencio Josué Rivera-Andrade
[…]
{affiliations}
J. Garrido, D. Silva-Muñiz, E. Riveiro, J. Rivera-Andrade
Department of Automation and Systems Engineering, Universidade de Vigo, Street Maxwell 9, Vigo, 36310, Pontevedra, Spain
Email Address: jgarri@uvigo.gal
J. Vales, P. López-Matencio
Department of Information Technologies and Communications, Technical University of Cartagena, Street Plaza del Hospital 1, Cartagena, 30202, Murcia, Spain
@@ 1 Introduction
Cable-driven parallel robots (CDPRs) are robotic systems where the end-effector’s position and orientation are controlled by adjusting the length of multiple cables, offering advantages in cost, performance, and adaptability [1]. Their structure, composed of cables and winches, enables larger workspaces and higher payload capacities compared to anthropomorphic robots [2]. Unlike other parallel robots like Delta or Stewart platforms, CDPRs are also more reconfigurable and capable of complex motions, making them ideal for applications that require extended workspaces and high load capacities [3, 4].
@@
However, their reliance on cables introduces a key limitation: cables can only operate under tension, not compression. As a result, stability depends on the specific setup of the systems. For instance, an eight-cable setup (four upper and four lower) provides precise control and robustness, while a four-cable configuration, though suitable for larger workspaces, is under-constrained and more prone to oscillations in the end-effector [5].
@@
Research has explored CDPR applications in diverse fields such as 3D printing [6], rehabilitation [7], storage and retrieval systems [8], sports tracking devices [9], precision agriculture [10] and load manipulation [11, 12]. Despite the aforementioned advantages, their industrial adoption remains limited [13, 14]. The main barriers to widespread implementation are the scarcity of commercial solutions, with only a few available, such as RBOT9 and CRANEBOT [15, 16], and challenges in meeting industrial safety and operational standards required for seamless integration with industrial controllers.
@@
Enhancing safety could significantly boost CDPR adoption, but research has mainly focused on cable breakage and collisions. This kind of solutions include adding passive auxiliary cables with clamping devices for redundancy [17] to reconfigurable architectures that actively relocate cable anchor points for task recovery after a cable failure [18], as well as collision-management techniques that address external impacts with humans or the environment [19], and internal self-interference [20].
@@
Besides, as stated, under-constrained CDPRs are more sensitive to external forces like wind, causing lateral disturbances that impact static positioning in pick-and-place tasks [21]. Accurate endpoint positioning is crucial, while trajectory errors during motion are more tolerable. Deviations of the end-effector due to external factors like wind or unintended operator contact are anomalies of the expected and safe behavior [22]. These anomalies can compromise grip quality or degrade outcomes in applications like laser inspection. Therefore, detecting these anomalies is the first step toward implementing actions to mitigate them.
@@
Although machine learning, especially deep learning, has improved anomaly detection [23, 24], many solutions are highly specialized and require extensive domain knowledge. Traditional methods like k-Nearest Neighbour, Decision Trees, or Discriminant Analysis have been applied to detect cable breakage or sensor failures in CDPRs [25], and Neural Networks have identified deviations from standard toolpaths [26]. Despite recent advances, CDPRs still lack practical anomaly detection strategies that can learn on the fly, without pre-labeled anomalies, and remain robust to wind gusts.
@@
We present an approach to detect unintended end-effector movements by analyzing anomalies in cable tensions without requiring additional external sensors or detailed environmental information. Our study focuses on conditions in which motion control maintains the robot’s pose while the motors remain powered and the motor brakes are switched off. These conditions occur, for instance, when the manipulator remains at rest while performing a picking or placing operation with its tool, as in typical pick-and-place tasks. We investigate whether disturbances from wind or other external forces can be detected solely through motor torque measurements based on servo amplifier current consumption. By doing so, any detected anomalies can prompt the postponement of subsequent operations until the anomaly is resolved. This approach enhances safety and provides redundancy for critical decisions. From an implementation perspective, cable tension are inferred directly from motor torques, avoiding the extra hardware, wiring, and calibration effort required by cable-mounted load cells as proposed in [27].
@@ 2 Static Stability of Under-Constrained Cable-Driven Parallel Robots
Stability describes the response of a mechanical system to a disturbance and is a fundamental property in the design and operation of CDPRs, particularly for under-constrained systems like our 4-cable experimental system (see Section 3), where gravity significantly influences performance [28]. A configuration is considered stable when a slight change in the cable forces does not permanently displace the platform or prevent it from returning to its original position once the disturbance disappears. Previous research has addressed the stability of CDPRs through analytical methods that involve the use of Lagrange multipliers and the computation of the Hessian matrix, as in [29] and [30], respectively. Specifically, following these works, stability is assured when the reduced Hessian ($\mathbf{H}_{\mathrm{r}}$) is positive definite.
@@
To simplify the model, we neglect the influence of pulley transmission on platform accelerations, cable tensions, and workspace boundaries. This assumption is supported by prior studies showing that omitting pulley effects reduces the workspace volume by approximately 1% when pulley radii are around 2.5% of the robot’s shortest edge [31]. Although pulley kinematics may temporarily amplify platform acceleration by up to 50% in worst-case scenarios [32], such transient effects are typically mitigated by the inner control loops of the drives and do not significantly impact the steady-state behavior under analysis.
@@
Using the notation introduced by [29], the static stability condition reads
@@
Under this static stability assumption, friction dissipates kinetic energy, leading the platform to a state of equilibrium over time. This result stems from the Lagrangian function $\mathcal{L}$, which relates the kinetic and potential energy to the system’s kinematics and dynamics. Following the criterion in [29], the reduced Hessian $\mathbf{H_{r}}$ of $\mathcal{L}$ must be positive definite for the equilibrium defined in Equation (2). Its expression is
@@ 3 Experimental setup
Inside the controller, kinematics, trajectory generation and control structures are implemented in Structured Text (IEC 61131-3 [33]) under the TwinCAT 3 runtime [34], with motor control compliant with PLCopen Motion Control standards [35]. Torque and position data is forwarded over EtherCAT to a digital I/O card and then to four EL7211 servo drives. Each servo drive powers an AM8112 servomotor that reels the cables onto the winch.
@@ 4.2 Model generation
Given the $C$ calibration points and setting the number of clusters to $K$, the GMM model parameters (means $\boldsymbol{\mu}_{k}$ and covariances $\boldsymbol{\Sigma}_{k}$, for $k$=$1,\ldots,K$) are fitted using the well-known Expectation-Maximization (EM) algorithm (e.g., see [36], Sect. 9.4.1) implemented in Python’s scikit-learn library through the GaussianMixture class. The number of cluster $K$ is selected as the one minimizing the Bayesian Information Criterion (BIC) for the model fit, which is commonly used in models aimed at anomaly detection (see [36], Sect. 9.4.3).
@@ 4.3 Anomaly detection
Once the model is fit, we can compute the distance from a new point $\boldsymbol{x}$ to the GMM model using the Mahalanobis distance. This is a common measure of the novelty of a point versus a (multi-)clustered data set, which we assume to exist since the conditions are deemed as statically stable during the calibration phase (as studied in Section 3.1). Similar approaches for the analysis of outliers in data series have been successfully applied in other domains [37, 38, 39]. In a multi-clustered framework the distance to the $k$-th cluster centroid is:
@@ 5 Experimental results
Figure 6 shows the results of the “method with GMM updates” on four additional experiments (#6, #7, #9, and #14). More specifically, Figure 6(a) shows the results of the experiment #6 in which the CDPR alternates between one-hour rest and one-hour steady wind periods. During each calm period (e.g., at 3 h, 5 h, and 9 h) the method is updated and corrects drift. Also, this figure shows how the algorithm detects all anomalies above the threshold (yellow bands) and switches back to normal classification almost as soon as the anomaly ends, with a TN rate of 99.4%. The few false positives (0.6% of the samples) appear during the wind-to-rest transition.
@@
Figure 6(b) shows the system under winds of varying intensity, as reflected by the variations in the relative distance signal. Although the signal fluctuates, it remains above the threshold, so the detector correctly identifies the whole period as anomalous. At about 3.5 h the wind stops, but the algorithm waits for a minimum settling time before it considers the system back to normal, so this brief interval is still labeled anomalous even the distance suggests that it could be a nominal state. Additionally, there is a false positive at around 4.25 h, probably caused by the alterations in the model calculation. For safety reasons, the system flags an anomaly, and once the model is correctly updated, it returns to identify the true negatives accurately.
@@
Figure 6(d) shows that the algorithm correctly identifies all anomalous areas. However, due to two recalibrations of the model at around 1.15 h, stabilisation does not occur until 2.2 h, temporarily delaying the accurate classification of the nominal states. This represents the most challenging scenario tested, yet the system still meets the required specifications. It successfully identifies all anomalous events and maintains an acceptable accuracy in detecting nominal conditions (Table 1). The delay in returning to normal classification is attributed to the temporary instability introduced by model adjustments.
As an example of anomaly detection, Figure 7 illustrates the motion of the load when an anomaly was detected in test file #14. This anomaly was caused by a maximum‐level (3) wind gust. Although the CDPR robot attempts to compensate for wind‐induced motion, a slight displacement of the payload remains perceptible (highlighted by the left and right arrows in the figure to facilitate identification of the movement). In other anomaly examples, the oscillatory movements are less visually evident due to the CDPR’s rapid compensation, yet they are still readily detected by our proposed algorithm through the resulting changes in the torques.

### v2

@@ Abstract
Keywords Anomaly detection  $\cdot$ Cable-Driven Parallel Robots  $\cdot$ Gaussian Mixture Models  $\cdot$ Robotics Control Systems  $\cdot$ Unsupervised Learning
@@ 1 Introduction
Cable-driven parallel robots (CDPRs) are robotic systems where the end-effector’s position and orientation are controlled by adjusting the length of multiple cables, offering advantages in cost, performance, and adaptability zarebidokiReview2022 . Their structure, composed of cables and winches, enables larger workspaces and higher payload capacities compared to anthropomorphic robots martin-parraNovel2024 . Unlike other parallel robots like Delta or Stewart platforms, CDPRs are also more reconfigurable and capable of complex motions, making them ideal for applications that require extended workspaces and high load capacities bobyMeasurement2021 ; izardCable2023 .
@@
However, their reliance on cables introduces a key limitation: cables can only operate under tension, not compression. As a result, stability depends on the specific setup of the systems. For instance, an eight-cable setup (four upper and four lower) provides precise control and robustness, while a four-cable configuration, though suitable for larger workspaces, is under-constrained and more prone to oscillations in the end-effector korayemDynamics2017 .
@@
Research has explored CDPR applications in diverse fields such as 3D printing qianCalibration2024 , rehabilitation huoDevelopment2024 , storage and retrieval systems bruckmannDevelopment2012 , sports tracking devices ghanatianExperimental2023a , precision agriculture prabhaCable2021 and load manipulation cullaFull2018 ; kimRemotely2022 . Despite the aforementioned advantages, their industrial adoption remains limited metillonPerformance2022 ; boumannDevelopment2019 . The main barriers to widespread implementation are the scarcity of commercial solutions, with only a few available, such as RBOT9 and CRANEBOT rbot9automationROCAP ; CRANEBOT , and challenges in meeting industrial safety and operational standards required for seamless integration with industrial controllers.
@@
Enhancing safety could significantly boost CDPR adoption, but research has mainly focused on cable breakage and collisions. This kind of solutions include adding passive auxiliary cables with clamping devices for redundancy caroFailure2020 to reconfigurable architectures that actively relocate cable anchor points for task recovery after a cable failure ramanFailure2022 , as well as collision-management techniques that address external impacts with humans or the environment rousseauHumancable2022 , and internal self-interference blanchetInterference2014 .
@@
Besides, as stated, under-constrained CDPRs are more sensitive to external forces like wind, causing lateral disturbances that impact static positioning in pick-and-place tasks husseinGeometric2018 . Accurate endpoint positioning is crucial, while trajectory errors during motion are more tolerable. Deviations of the end-effector due to external factors like wind or unintended operator contact are anomalies of the expected and safe behavior taghaviDevelopment2018 . These anomalies can compromise grip quality or degrade outcomes in applications like laser inspection. Therefore, detecting these anomalies is the first step toward implementing actions to mitigate them.
@@
Although machine learning, especially deep learning, has improved anomaly detection choiDeep2021 ; pangDeep2022 , many solutions are highly specialized and require extensive domain knowledge. Traditional methods like k-Nearest Neighbour, Decision Trees, or Discriminant Analysis have been applied to detect cable breakage or sensor failures in CDPRs bettegaLoad2024 , and Neural Networks have identified deviations from standard toolpaths jabbariaslAdaptive2017a . Despite recent advances, CDPRs still lack practical anomaly detection strategies that can learn on the fly, without pre-labeled anomalies, and remain robust to wind gusts.
@@
We present an approach to detect unintended end-effector movements by analyzing anomalies in cable tensions without requiring additional external sensors or detailed environmental information. Our study focuses on conditions in which motion control maintains the robot’s pose while the motors remain powered and the motor brakes are switched off. These conditions occur, for instance, when the manipulator remains at rest while performing a picking or placing operation with its tool, as in typical pick-and-place tasks. We investigate whether disturbances from wind or other external forces can be detected solely through motor torque measurements based on servo amplifier current consumption. By doing so, any detected anomalies can prompt the postponement of subsequent operations until the anomaly is resolved. This approach enhances safety and provides redundancy for critical decisions. From an implementation perspective, cable tension are inferred directly from motor torques, avoiding the extra hardware, wiring, and calibration effort required by cable-mounted load cells as proposed in gaoDetection2024 .
@@ 2 Static Stability of Under-Constrained Cable-Driven Parallel Robots
Stability describes the response of a mechanical system to a disturbance and is a fundamental property in the design and operation of CDPRs, particularly for under-constrained systems like our 4-cable experimental system (see Section 3), where gravity significantly influences performance pottCableDriven2018 . A configuration is considered stable when a slight change in the cable forces does not permanently displace the platform or prevent it from returning to its original position once the disturbance disappears. Previous research has addressed the stability of CDPRs through analytical methods that involve the use of Lagrange multipliers and the computation of the Hessian matrix, as in carricatoGeometricoStatic2010 and abbasnejadDirect2015 , respectively. Specifically, following these works, stability is assured when the reduced Hessian ($\mathbf{H}_{\mathrm{r}}$) is positive definite.
@@
To simplify the model, we neglect the influence of pulley transmission on platform accelerations, cable tensions, and workspace boundaries. This assumption is supported by prior studies showing that omitting pulley effects reduces the workspace volume by approximately 1% when pulley radii are around 2.5% of the robot’s shortest edge pottInfluence2012 . Although pulley kinematics may temporarily amplify platform acceleration by up to 50% in worst-case scenarios idaRestRest2019 , such transient effects are typically mitigated by the inner control loops of the drives and do not significantly impact the steady-state behavior under analysis.
@@
Using the notation introduced by carricatoGeometricoStatic2010 , the static stability condition reads
@@
Under this static stability assumption, friction dissipates kinetic energy, leading the platform to a state of equilibrium over time. This result stems from the Lagrangian function $\mathcal{L}$, which relates the kinetic and potential energy to the system’s kinematics and dynamics. Following the criterion in carricatoGeometricoStatic2010 , the reduced Hessian $\mathbf{H_{r}}$ of $\mathcal{L}$ must be positive definite for the equilibrium defined in Equation (2). Its expression is
@@ 3 Experimental setup
Inside the controller, kinematics, trajectory generation and control structures are implemented in Structured Text (IEC 61131-3 iecProgrammable2013 ) under the TwinCAT 3 runtime beckhoffautomationgmbh&co.kgTwinCAT , with motor control compliant with PLCopen Motion Control standards plcopenFunction2008 . Torque and position data is forwarded over EtherCAT to a digital I/O card and then to four EL7211 servo drives. Each servo drive powers an AM8112 servomotor that reels the cables onto the winch.
@@ 4.2 Model generation
Given the $C$ calibration points and setting the number of clusters to $K$, the GMM model parameters (means $\boldsymbol{\mu}_{k}$ and covariances $\boldsymbol{\Sigma}_{k}$, for $k$=$1,\ldots,K$) are fitted using the well-known Expectation-Maximization (EM) algorithm (e.g., see bishopPattern2006 , Sect. 9.4.1) implemented in Python’s scikit-learn library through the GaussianMixture class. The number of cluster $K$ is selected as the one minimizing the Bayesian Information Criterion (BIC) for the model fit, which is commonly used in models aimed at anomaly detection (see bishopPattern2006 , Sect. 9.4.3).
@@ 4.3 Anomaly detection
Once the model is fit, we can compute the distance from a new point $\boldsymbol{x}$ to the GMM model using the Mahalanobis distance. This is a common measure of the novelty of a point versus a (multi-)clustered data set, which we assume to exist since the conditions are deemed as statically stable during the calibration phase (as studied in Section 3.1). Similar approaches for the analysis of outliers in data series have been successfully applied in other domains vales-alonsoNonsupervised2023a ; zhengElectromyographicBased2025 ; belkhoucheRobust2022 . In a multi-clustered framework the distance to the $k$-th cluster centroid is:
@@ 5 Experimental results
Figure 6 shows the results of the “method with GMM updates” on four additional experiments (#6, #7, #9, and #14). More specifically:
@@
  • Figure 6(a) shows the results of the experiment #6 in which the CDPR alternates between one-hour rest and one-hour steady wind periods. During each calm period (e.g., at 3 h, 5 h, and 9 h) the method is updated and corrects drift. Also, this figure shows how the algorithm detects all anomalies above the threshold (yellow bands) and switches back to normal classification almost as soon as the anomaly ends, with a TN rate of 99.4%. The few false positives (0.6% of the samples) appear during the wind-to-rest transition.
  • Figure 6(b) shows the system under winds of varying intensity, as reflected by the variations in the relative distance signal. Although the signal fluctuates, it remains above the threshold, so the detector correctly identifies the whole period as anomalous. At about 3.5 h the wind stops, but the algorithm waits for a minimum settling time before it considers the system back to normal, so this brief interval is still labeled anomalous even when the distance suggests that it could be a nominal state. Additionally, there is a false positive at around 4.25 h, probably caused by the alterations in the model calculation. For safety reasons, the system flags an anomaly, and once the model is correctly updated, it returns to identify the true negatives accurately.
@@
  • In the test shown in Figure 6(d), the algorithm perfectly identifies all outlier regions. However, because stability problems arise around 1.5 h, the model does not reach steady state until 1.64 h, which slightly delays accurate classification of the nominal states. This represents the most challenging scenario tested, yet the system still meets the required specifications. It successfully identifies all anomalous events and maintains an acceptable accuracy in detecting nominal conditions (Table 1). The delay in returning to normal classification is attributed to the temporary instability introduced by external winds, which delays model adjustments.
As an example of anomaly detection, Figure 7 illustrates the motion of the load when an anomaly was detected in test file #14. This anomaly was caused by a maximum-level (3) wind gust. Although the CDPR robot attempts to compensate for wind-induced motion, a slight displacement of the payload remains perceptible (highlighted by the left and right arrows in the figure to facilitate identification of the movement). In other anomaly examples, the oscillatory movements are less visually evident due to the CDPR’s rapid compensation, yet they are still readily detected by our proposed algorithm through the resulting changes in the torques.